### 12 Mar 2026 v1

# InSpatio-WorldFM: An Open-Source Real-Time Generative Frame Model

InSpatio-WorldFM: An Open-Source Real-Time Generative Frame Model
Xiaoyu Zhang, Weihong Pan, Zhichao Ye, Jialin Liu, Yipeng Chen, Nan Wang,
Xiaojun Xiang, Weijian Xie, Yifu Wang, Haoyu Ji, Siji Pan, Zhewen Le, Jing Guo,
Xianbin Liu, Donghui Shen, Ziqiang Zhao, Haomin Liu, Guofeng Zhang
InSpatio Team
We present InSpatio-WorldFM, an open-source real-time frame model for spatial intelligence. Unlike video-based world models that rely on sequential frame generation and incur substantial latency due to window-level processing, InSpatio-WorldFM adopts a frame-based paradigm that generates each frame independently, enabling low-latency real-time spatial inference. By enforcing multi-view spatial consistency through explicit 3D anchors and implicit spatial memory, the model preserves global scene geometry while maintaining fine-grained visual details across viewpoint changes. We further introduce a progressive three-stage training pipeline that transforms a pretrained image diffusion model into a controllable frame model and finally into a real-time generator through few-step distillation. Experimental results show that InSpatio-WorldFM achieves strong multi-view consistency while supporting interactive exploration on consumer-grade GPUs, providing an efficient alternative to traditional video-based world models for real-time world simulation. Project Website: https://inspatio.github.io/worldfm/ Github: https://github.com/inspatio/worldfm
Figure 1: Examples of generated worlds across diverse styles, including photorealistic, science-fiction, game-like, and artistic environments. The joystick interface enables real-time interactive exploration with negligible latency.
1 Introduction
Recent advances in generative models (Rombach et al., 2022; Labs, 2024; Labs et al., 2025; Team, 2024; Blattmann et al., 2023; Peebles and Xie, 2023) have significantly improved visual realism and temporal coherence. These models learn rich motion patterns and camera dynamics from large-scale video datasets Ling et al. (2024); Zhou et al. (2018); Yao et al. (2020), which has sparked growing interest in using generative models to simulate persistent realistic environments. As a result, the research focus is gradually shifting from traditional text-to-video generation (Bar-Tal et al., 2024; Singer et al., 2022; Brooks et al., 2024a; Yang et al., 2024; Brooks et al., 2024b; Team, 2025) toward the broader goal of building world models capable of representing and interacting with structured environments (Zhou et al., 2025b; Yang et al., 2025; Chung et al., 2023; Yu et al., 2024, 2025; Gao et al., 2024), such as HY-World (Team et al., 2025) and LingBot-World (Team et al., 2026) , as well as the models that incorporate predefined rules or constraints (Liu et al., 2024; Zhang et al., 2024b; Raistrick et al., 2024), such as Genie 3 (DeepMind, 2025).
Despite this progress, most existing world models (Huang et al., 2025; Wu et al., 2025; Ren et al., 2025; Ding et al., 2025; Zhang et al., 2025; Sun et al., 2025) are still built upon video generation architectures. Leveraging video models allows these systems to inherit strong priors about motion, appearance, and camera movement learned from large-scale video data. However, this paradigm also introduces several fundamental limitations.
Interactive latency remains inevitable. Most video-based world models generate frames sequentially within a temporal window. With bidirectional attention and full-window decoding, each generation step must process all frames in the window, resulting in substantial inference overhead. Even with distillation-based acceleration, this window-level dependency fundamentally limits the ability of such models to support truly real-time interaction.
Spatial errors accumulate over time. Video models are primarily optimized for short-term temporal continuity rather than long-term spatial consistency. As generation progresses, small spatial inaccuracies may accumulate, leading to structural drift or inconsistencies in scene geometry. Without explicit mechanisms to enforce global spatial constraints, correcting these accumulated errors remains challenging.
Recently, World Labs proposed a real-time frame model (RTFM) (WorldLabs, 2025), exploring the potential of frame-based world modeling. However, the technical details provided are quite limited and the source code is not publicly available. To further explore this direction and address the limitations of video-based world models, we propose InSpatio-WorldFM, a frame-based world model designed for real-time spatial reasoning and generation. Unlike video-based approaches that model the world as a sequence of dependent frames, InSpatio-WorldFM directly incorporates spatial structure into the generation of individual frames. By conditioning each frame on explicit spatial information, the model can maintain consistent scene geometry while enabling low-latency frame synthesis. This design allows the system to generate spatially coherent observations while supporting real-time interaction. We introduce several key components:
  • Multi-view consistent training data curation. We construct training data with explicit multi-view consistency, allowing the model to learn stable spatial relationships across viewpoints.
  • Progressive three-stage training pipeline that evolves a foundation image generator (Stage I), into a controllable frame model with spatial memory (Stage II), and finally into a real-time few-step generator (Stage III).
  • Real-time generation enabled by few-step distillation (Yin et al., 2024) through 2-step denoising. Experimental results demonstrate that InSpatio-WorldFM can generate frames with strong multi-view consistency while achieving real-time inference without interactive latency.
Together, these components enable InSpatio-WorldFM to serve as an efficient foundation for real-time spatial intelligence, providing a practical alternative to traditional video-based world models.
Figure 2: Overview. In the offline stage, a multi-view-consistent model generates plausible observations that provide 3D anchors and reference appearances. In the online stage, frame model performs fast real-time inference while updating scene content at keyframes.
2 InSpatio-WorldFM
2.1 Overview
As illustrated in Fig. 2, the framework of InSpatio-WorldFM consists of an offline stage and an online stage. In the offline stage, a single image is used as input to a multi-view consistent model (e.g., (Zhou et al., 2025a; Gao et al., 2024)) to generate multi-view consistent observations, which provide reference appearances, while the 3D anchors can be obtained using a reconstruction model (e.g.,  (Wang et al., 2025, 2024b, 2024a)). Alternatively, a simpler approach is to generate a panoramic image, which can also provide 360° scene constraints, and several open-source methods are already available for this task (Team et al., 2025; Feng et al., 2023; Zhang et al., 2024a). In the online stage, we introduce an efficient frame model that performs real-time inference and generation.
2.2 Formulation
We formulate InSpatio-WorldFM as a conditional generative frame model that synthesizes novel-view images of a 3D scene from a single reference image under user-defined camera motion. Let $x_{\text{ref}}\in\mathbb{R}^{H\times W\times 3}$ denotes the reference image with its associated camera pose $\pi_{\text{ref}}=(K_{\text{ref}},E_{\text{ref}})$, where $K$ and $E$ represent the camera intrinsic and extrinsic matrices, respectively. Given a target camera pose $\pi_{\text{tgt}}$, our goal is to generate the corresponding target-view image $x_{\text{tgt}}$ that is geometrically consistent with $x_{\text{ref}}$.
We approach this task within a latent diffusion framework. Let $\mathcal{E}$ and $\mathcal{D}$ denote the encoder and decoder of a pretrained variational autoencoder (VAE), respectively. The generation is formulated as learning a conditional denoising model $\epsilon_{\theta}$ that reverses a forward diffusion process in the latent space $z=\mathcal{E}(x)$. Specifically, the model learns to predict the noise $\epsilon$ added to the target latent $z_{\text{tgt}}$ at diffusion timestep $t$:
$\mathcal{L}=\mathbb{E}_{z_{\text{tgt}},\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathcal{C})\right\|^{2}\right],$
where $z_{t}=\alpha_{t}z_{\text{tgt}}+\sigma_{t}\epsilon$ is the noised latent at timestep $t$ following a predefined noise schedule $(\alpha_{t},\sigma_{t})$, and $\mathcal{C}=\{x_{\text{ref}},\pi_{\text{ref}},\pi_{\text{tgt}},\hat{x}_{\text{tgt}}\}$ represents the full condition set. Here $\hat{x}_{\text{tgt}}$ denotes the point cloud rendering at the target viewpoint with 3D foundation models (Wang et al., 2025, 2024b, 2024a; Li and Snavely, 2018), which serves as an explicit 3D spatial anchor.
To achieve real-time / interactive rendering, we adopt a three-stage training pipeline that progressively evolves a foundation image generator into an efficient real-time frame model:
  • Stage I: Pre-Training. We establish a high-fidelity image generation prior by selecting an efficient and expressive Diffusion Transformer as our backbone. The choice of the base model is driven by both generation quality and computational efficiency, as the latter directly impacts real-time deployment feasibility.
  • Stage II: Middle-Training. We transform the pretrained image generator into a controllable frame model with spatial memory. This stage involves constructing dedicated training data from real-world video datasets and synthetic environments, introducing architectural modifications for camera-conditioned generation, and developing a hybrid spatial memory mechanism that combines explicit 3D anchors with implicit neural memory.
  • Stage III: Post-Training. We distill the multi-step diffusion model into an efficient few-step generator through distribution matching distillation, enabling real-time / interactive rendering on consumer-grade GPUs.
2.3 Pre-Training
The choice of the base image generation model is critical, as it must simultaneously satisfy two requirements: high-fidelity generation with physically plausible textures and scene geometry, and computational efficiency sufficient for downstream real-time deployment. We select PixArt-$\Sigma$ (Chen et al., 2024) as our foundation model. PixArt-$\Sigma$ is an efficient Diffusion Transformer (DiT) (Peebles and Xie, 2023) for text-to-image generation that achieves quality competitive with state-of-the-art models at significantly lower computational cost, providing an effective balance between generation quality and inference throughput that is well-suited as the backbone for a real-time frame model.
2.4 Middle-Training
The middle-training stage transforms the pretrained image diffusion model into a controllable frame model capable of generating spatially consistent and interactive visual content. Although foundation diffusion model demonstrates strong potential in high-fidelity image synthesis, it is inherently limited to generating single, isolated images without any notion of spatial coherence or interactive controllability. We therefore construct dedicated training data and introduce architectural modifications to endow the model with action-conditioned control and spatial consistency.
2.4.1 Fundamental Frame Model
To achieve true real-time performance, we adopt a minimal frame-based architecture as the online inference model, generating each frame independently. To maintain multi-view consistency across successive inferences without relying on multi-frame processing, we integrate explicit 3D anchors (point cloud rendering) with implicit spatial memory (reference frame attention), enabling the model to preserve geometric coherence while operating on individual frames.
Figure 3: The pipeline of InSpatio-WorldFM. The left part illustrates the conditional novel-view synthesis pipeline of WorldFM. WorldFM takes a reference image $x_{\text{ref}}$ (implicit scene memory), noisy latents $z_{t}$, and point cloud rendering $\hat{x}_{\text{tgt}}$ (explicit 3D anchor) as inputs, which are spatially concatenated along the width dimension. Reference pose $\pi{\text{ref}}$ and target pose $\pi_{\text{tgt}}$ are also injected as control signals. The frame-based Diffusion Transformer Blocks process these conditions and synthesizes the target view $I_{\text{tgt}}$ in real-time via Distribution Matching Distillation (DMD). The right part shows the detailed architecture of the DiT blocks in WorldFM. Camera geometry control is achieved through the Projection Relative Position Embedding (PRoPE) strategy, enhancing cross-view geometric reasoning. The hybrid spatial memory mechanism combines point cloud rendering (explicit 3D anchor) and reference image (implicit memory), interacting solely through self-attention to achieve robust 3D consistency.
Model Architecture. InSpatio-WorldFM generates target-view images by receiving a reference image, noisy target latents, and user-defined camera motion as inputs.
We adopt a self-attention-only transformer architecture for condition injection. Empirically, we find that injecting condition information through self-attention yields higher generation quality than cross-attention alternatives. The input to the transformer is formed by spatially concatenating three components along the width dimension: the noisy target latent $z_{t}$, the point cloud rendering of the target viewpoint (cond1), and the reference frame image (cond2). All three components are independently projected into patch tokens via a shared patch embedding layer and augmented with sinusoidal positional embeddings. After passing through the full transformer, the output is split along the width dimension and only the target portion is retained as the final prediction.
Camera Pose Encoding. Encoding camera geometry into the transformer is essential for controllable novel-view generation. We explore three strategies for injecting camera pose information:
  • Plücker ray embedding (He et al., 2025; Agarwal et al., 2025; Alhaija et al., 2025). For each patch token, the 6-dimensional Plücker coordinates $(\mathbf{o}\times\mathbf{d},\mathbf{d})$ are computed in world coordinates, where $\mathbf{o}$ is the camera origin and $\mathbf{d}$ is the ray direction through the patch center. These features are projected to the hidden dimension via a two-layer MLP and added to the patch embeddings. This approach provides an explicit per-token geometric prior but encodes camera information additively without directly modulating the attention computation.
  • Projected Relative Positional Encoding (PRoPE) (Li et al., 2025). PRoPE integrates camera geometry directly into the attention mechanism by applying camera-dependent linear transformations to the query, key, and value tensors. Specifically, given camera projection matrices $P_{i}$ for each view $i$, PRoPE applies $P_{i}^{\top}$ to queries and $P_{i}^{-1}$ to keys/values, combined with 2D rotary position embeddings for intra-image spatial structure. This formulation enables the attention mechanism to reason about cross-view geometric correspondences natively.
  • Pure parametric injection (Bai et al., 2025). Camera pose parameters (rotation and translation matrices) are directly mapped to token-level embeddings through a learned MLP and added to the hidden representations, without imposing explicit geometric structure such as ray directions or projection matrices.
We adopt PRoPE for InSpatio-WorldFM, as it exhibits the fastest convergence and the most stable camera control among the three strategies in our experiments. Compared with Plücker ray embedding, which encodes camera information additively, PRoPE directly modulates the attention computation through camera-dependent linear transformations, enabling the attention mechanism to reason about cross-view geometric correspondences natively.
Hybrid Spatial Memory. A core challenge in maintaining spatial consistency across generated frames is preserving a coherent memory of the 3D scene. RTFM (WorldLabs, 2025) relies on posed frames as primitive spatial memory, while StarGen (Zhai et al., 2025) warps features extracted from posed keyframes to provide spatial conditioning. To achieve more persistent and stable spatial memory, we adopt a hybrid design that integrates explicit 3D anchors with implicit neural memory:
  • Explicit anchors. For each target viewpoint, a point cloud rendering and the corresponding target camera pose provide stable 3D geometric constraints, serving as global 3D priors that anchor the generation in 3D space. While point clouds effectively maintain coarse geometric structures, they are complemented by implicit memory to preserve fine-grained appearance details.
  • Implicit memory. The reference frame and its associated camera pose supply appearance information from a previously observed viewpoint. The transformer attends to these tokens through the self-attention mechanism, allowing the model to implicitly retrieve and transfer relevant visual content to the target view.
By combining both forms of spatial memory, InSpatio-WorldFM achieves robust consistency: explicit anchors maintain coarse geometric structures and provide global 3D priors, while implicit memory preserves fine-grained appearance details and enables the model to hallucinate plausible content in unobserved regions.
Training Data. We construct training pairs from three sources: publicly available videos (e.g., internet videos and video datasets such as DL3DV (Ling et al., 2024), RealEstate10K (Zhou et al., 2018)), our own captured videos, and the synthetic data generated using Unreal Engine (UE) (Epic Games, 2023). For each real video clip, we randomly sample 16 frames and apply a feedforward reconstruction model (e.g., MapAnything (Keetha et al., 2025)) to estimate per-frame camera poses and depth maps. From these 16 frames, 4 are selected as the reference frame group to construct the global point cloud, while the remaining 12 serve as training targets. For each target frame, the reference frame is selected as the temporally closest one from the 4-frame reference group. The point cloud rendering at the target viewpoint is obtained by projecting the global point cloud onto the target camera plane. To ensure robustness, we apply random shuffling and masking strategies, simulating the disorder and discreteness encountered in real-world scenarios.
Training Strategy. We introduce several training strategies to encourage stable geometric learning and robust spatial memory:
  • Noise schedule biasing. To prioritize learning stable geometric structure over fine-grained texture details, we increase the sampling probability of high-noise timesteps during training. This biased schedule encourages the model to first capture the coarse spatial layout before refining appearance.
  • Progressive condition injection. We observe that simultaneously exposing the model to both the explicit anchor (point cloud rendering) and the implicit memory (reference frame) from the beginning of training leads to rapid overfitting to the explicit anchor, as it provides a much stronger and more direct signal. This causes the model to neglect the implicit memory pathway. To ensure robustness and avoid overfitting, we adopt a progressive injection strategy: in the early phase of training, only the reference frame (implicit memory) is provided as a condition, forcing the model to develop a robust ability to extract and utilize implicit spatial cues. The explicit anchor is then gradually introduced to enhance control precision.
  • Random anchor masking. In the later stages of training, we randomly mask the explicit anchor with a certain probability. This regularization encourages the model to maintain its ability to leverage implicit memory, preventing it from becoming overly dependent on the explicit 3D prior.
2.4.2 Finetuning with Synthetic Data
Although the fundamental frame model trained on real data acquires reasonable and geometrically consistent generation capabilities, the depth and pose estimates produced by feedforward reconstruction models inevitably contain errors. These inaccuracies introduce inter-view inconsistencies in the point cloud renderings, which undermine the stability of viewpoint transitions and content persistence.
To address this, we construct a synthetic dataset using Unreal Engine (Epic Games, 2023) that provides precise ground-truth camera poses and depth information. Specifically, our approach begins by selecting an initial camera pose from semantically valid regions within the scene. The trajectory is then synthesized using either stochastic motion sampling or pre-defined motion templates. Throughout generation, we enforce spatial constraints through collision avoidance to guarantee viewpoint validity. Training pairs are then constructed analogously to the real data pipeline: 4 frames serve as the reference group and 12 as targets.
We finetune the fundamental frame model on this synthetic dataset for a limited number of steps. The controlled exposure to synthetic data is intentional — excessive finetuning on synthetic distributions would compromise the model’s ability to generate realistic appearances on natural images. Empirically, we find that even a small amount of synthetic finetuning yields a significant improvement in the stability of camera-controlled viewpoint transitions, suggesting that the model effectively leverages the geometric precision of synthetic data to refine its spatial reasoning without losing its natural appearance priors.
Figure 4: Qualitative results of teacher model.
Figure 5: Qualitative results of InSpatio-WorldFM.
Figure 6: Qualitative results of InSpatio-WorldFM.
Figure 7: Qualitative results of InSpatio-WorldFM.
Figure 8: Qualitative results of InSpatio-WorldFM. Across observations at varying viewing distances, the content and fine details generated by the FrameModel remain consistent.
2.5 Post-Training
The goal of post-training is to transform the multi-step frame model from Stage II into an efficient generator capable of real-time interactive rendering. We achieve this through Distribution Matching Distillation (DMD) (Yin et al., 2024), which distills the multi-step diffusion model into a few-step generator with minimal loss in spatial consistency and visual fidelity.
Distribution Matching Distillation. DMD trains a few-step diffusion student to match the output distribution of a pretrained multi-step diffusion model. The core idea is to minimize an approximate KL divergence between the real distribution (defined by the base model) and the synthetic distribution (produced by the generator). This is achieved by maintaining two diffusion models — a frozen copy of the base model estimating the real score, and a dynamically-updated model continually trained on generator outputs estimating the fake score — and using the difference between their denoising predictions as the gradient signal to update the generator. A complementary regression loss on pre-computed noise-image pairs from the base model’s deterministic sampler stabilizes training and preserves mode coverage. Extending the idea of Variational Score Distillation (VSD), DMD enables few-step inference acceleration while maintaining generation quality.
Adaptation for InSpatio-WorldFM. We apply DMD to distill our middle-trained frame model into a few-step generator. Through extensive experimentation, we arrive at two key findings:
  • Two-step denoising outperforms one-step. While DMD was originally designed for one-step generation, we find that a 2-step denoising schedule produces noticeably sharper details and better preserves fine-grained textures. Single-step denoising from pure noise can recover coarse geometric structure effectively but struggles to reconstruct fine-grained details in a single forward pass. The additional denoising step provides a dedicated refinement phase that addresses this limitation.
  • Intermediate timestep selection is critical. In the 2-step schedule, the first step denoises from $t=T$ to an intermediate timestep $t_{\text{mid}}$, and the second step denoises from $t_{\text{mid}}$ to $t=0$. If $t_{\text{mid}}$ is set too large (i.e., the first step terminates at a still high-noise state), the second step must perform single-step denoising from a high noise level — encountering the same difficulty as 1-step generation in recovering fine details. Through systematic evaluation, we find that $t_{\text{mid}}=200$ (on a 1000-step schedule) yields the best balance: the first step handles the majority of denoising and establishes the coarse spatial structure, while the second step refines from a relatively clean state where fine detail recovery is effective.
The distilled model maintains spatial consistency and visual fidelity over extended interactive sessions, producing temporally coherent frames without noticeable drift even during prolonged exploration.
3 Evaluation
We assess the spatial consistency and generation quality of InSpatio-WorldFM by examining the visual outputs from both the middle-trained fundamental frame model and the post-trained distilled variant across diverse scene types.
As illustrated in Fig. 4, each example consists of a reference image alongside a sequence of 10 frames rendered from different camera viewpoints. The fundamental frame model exhibits strong multi-view coherence, maintaining geometric structure and appearance consistency across substantial viewpoint changes. The generated frames preserve fine-grained details while adapting naturally to novel perspectives, demonstrating the model’s capacity to reason about 3D spatial relationships.
To enable real-time interaction, we apply distribution matching distillation to obtain a distilled model for low-latency deployment. In its baseline configuration, InSpatio-WorldFM runs at approximately 10 FPS at 512×512 resolution on a single NVIDIA A100 GPU with around 50-70 ms interaction latency. Owing to its low GPU memory consumption, it also enables interactive exploration on consumer-grade GPUs and achieves 7 fps on RTX 4090 with single-step inference. With engineering optimizations such as KV-cache management and efficient VAE latent caching, the inference speed can be further significantly improved. While distillation introduces a quality-speed trade-off, the perceptual difference remains minimal in practice. Visual comparison between the fundamental model and the distilled variant reveals that InSpatio-WorldFM successfully retains spatial coherence and geometric accuracy. The accelerated model continues to generate plausible content across diverse camera trajectories without noticeable degradation in structural consistency or the emergence of visual artifacts, confirming that our distillation approach effectively balances real-time performance with generation fidelity.
In long rendering tests under different historical frame contexts, InSpatio-WorldFM maintains consistent geometry and fine details across varying viewing distances, demonstrating the effectiveness of the memory strategy in guiding generation.
4 Discussion and Conclusions
In this report, we introduce InSpatio-WorldFM, an open-source real-time generative frame model designed for spatial intelligence. By adopting a frame-based generation paradigm and incorporating explicit 3D anchors with implicit spatial memory, the model maintains multi-view spatial consistency while enabling low-latency real-time inference.
4.1 Limitations
Despite the progress made, achieving an immersive real-time generative world model still faces several challenges.
  • Stable generation of dynamic content. Both the frame-based model and the multi-view consistency training data contain limited dynamic content, which makes it difficult for the model to generate dynamic scenes with high quality and stability.
  • Limited motion boundary. The current historical memory relies on multi-view consistent observations or panoramic observations. However, these generation models suffer from high computational complexity and substantial memory consumption, restricting them to offline operation, which inevitably introduces motion boundaries during online inference.
  • Interactive visual stability The frame-based modeling strategy reduces interaction latency and improves responsiveness. However, due to the lack of temporal constraints between consecutive frames, the visual stability is limited, leading to noticeable frame jitter during interaction.
4.2 Future Work
Real-time spatial inference still has significant room for improvement. The frame-based inference architecture can benefit from many mature acceleration techniques, such as linear attention, efficient caching mechanisms, and various VAE optimizations, to further reduce computational cost and improve FPS. These improvements would enable spatial inference to be performed more efficiently on edge devices. Using Gaussian Splatting (GS) primitives as 3D anchors could further enhance visual fidelity and reflective effects. Beyond efficiency and visual fidelity, we will focus on two key directions: (1) improving the generation of dynamic content, and (2) real-time expansion of the generation range. These efforts will provide the foundation for efficient real-world modeling and spatial inference at effectively unbounded scales.
5 Contributors
Model: Xiaoyu Zhang, Weihong Pan, Zhichao Ye, Yipeng Chen
Data: Jialin Liu, Xiaojun Xiang, Weijian Xie
Engineering: Nan Wang, Siji Pan, Jing Guo, Xianbin Liu, Ziqiang Zhao
Applications: Haomin Liu, Haoyu Ji, Donghui Shen
Art Designer: Yifu Wang, Zhewen Le
Project Lead: Guofeng Zhang, Xiaoyu Zhang
Acknowledgments
We thank Zhenzhou Fan, Jingfan Huang, Zhiwei Huang, Jingbo Liu, Chaoran Tian, Shao Zhao (listed alphabetically by last name) for their valuable discussions and assistance.